# Functional Graph Convolutional Networks: A unified multi-task and multi-modal learning framework to facilitate health and social-care insights

### v1

@@ Appendix C Additional tables
variable short description   $*{}^{*}$ (l) quality of life index max $*{}^{*}$ (l) maximum of grip strength measure recall $*{}^{*}$ (l) number of words recalled in the first trial $*{}^{*}$ (l) body mass index $*{}^{*}$ (l) activities of daily living index $*{}^{*}$ (l) sum of five daily activities $*{}^{*}$ (l) large muscle index $*{}^{*}$ (l) mobility index $*{}^{*}$ (l) grossmotor skills index $*{}^{*}$ (l) finemotor skills index income $*{}^{*}$ (l) household able to make ends meet $*{}^{*}$ (l) depression index with EURO-D scale income $*{}^{*}$ (l) household income percentiles heart attack (l) 1 if the subject ever had the disease, 0 o.w. high cholesterol (l) 1 if the subject ever had the disease, 0 o.w. stroke (l) 1 if the subject ever had the disease, 0 o.w. diabetes (l) 1 if the subject ever had the disease, 0 o.w. COPD (l) 1 if the subject ever had the disease, 0 o.w. cancer (l) 1 if the subject ever had the disease, 0 o.w. ulcer (l) 1 if the subject ever had the disease, 0 o.w. parkinson (l) 1 if the subject ever had the disease, 0 o.w. cataracts (l) 1 if the subject ever had the disease, 0 o.w. hip fracture (l) 1 if the subject ever had the disease, 0 o.w. age (s) age of the subject education years (s) years of education n children (a) number of children that are still alive numeracy (a) mathematical performance drinking behav (a) times a patient drunk in the last 6 months hospital days (a) days spent at the hospital in the last 6 months hospitalizations (a) number of hospitalizations in the last 6 months GP $*{}^{*}$ (c) number of doctor visits within the study: high, medium, low $*{}^{*}$ (c) female or male $*{}^{*}$ (c) being vaccinated during childhood ever $*{}^{*}$ (c) ever smoked daily

### v2

@@ Appendix C Additional tables
variable short description   CASP (l) quality of life index max grip (l) maximum of grip strength measure recall test (l) number of words recalled in the first trial bmi (l) body mass index adlwa (l) activities of daily living index adla (l) sum of five daily activities lgmuscle (l) large muscle index mobilityind (l) mobility index grossmotor (l) grossmotor skills index finemotor (l) finemotor skills index income perception (l) household able to make ends meet eurod (l) depression index with EURO-D scale income percentile (l) household income percentiles heart attack (l) 1 if the subject ever had the disease, 0 o.w. high cholesterol (l) 1 if the subject ever had the disease, 0 o.w. stroke (l) 1 if the subject ever had the disease, 0 o.w. diabetes (l) 1 if the subject ever had the disease, 0 o.w. COPD (l) 1 if the subject ever had the disease, 0 o.w. cancer (l) 1 if the subject ever had the disease, 0 o.w. ulcer (l) 1 if the subject ever had the disease, 0 o.w. parkinson (l) 1 if the subject ever had the disease, 0 o.w. cataracts (l) 1 if the subject ever had the disease, 0 o.w. hip fracture (l) 1 if the subject ever had the disease, 0 o.w. age (s) age of the subject education years (s) years of education n children (a) number of children that are still alive numeracy (a) mathematical performance drinking behav (a) times a patient drunk in the last 6 months hospital days (a) days spent at the hospital in the last 6 months hospitalizations (a) number of hospitalizations in the last 6 months GP visits (c) number of doctor visits within the study: high, medium, low gender (c) female or male vaccinated (c) being vaccinated during childhood ever smoked (c) ever smoked daily